Two parks are renting camping sites. Camp happy face rents their camp site for $250 base free and $30 per night. Camp doodleface rents their campsite for $50 per night with a $50 base free. Which camp has the best deal if you want to rent it for 12 nights?
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

250 + 30x = 50 + 50x
200 = 20x
x = 10 At 10 nights they both cost the same.
At 12 nights Happy Face is cheaper
Proof:
250 + 30(12) = 610
50 + 50(12) = 650

Answer by MathTherapy(10552)


Two parks are renting camping sites. Camp happy face rents their camp site for $250 base free and $30 per night. Camp doodleface rents their campsite for $50 per night with a $50 base free. Which camp has the best deal if you want to rent it for 12 nights?
Camp Happy Face's cost, based on N nights: 30N + 250
Camp Doodleface's cost, based on N nights: 50N + 50
Camp Happy Face's cost, for 12 nights: 30(12) + 250 = $610
Camp Doodleface's cost, for 12 nights: 50(12) + 50 = $650
Can you now choose the cheaper of the 2?